The radius and height of a cylinder are in the ratio 5:7 and its volume is 550cm^3. Find its radius and height.
1
Expert's answer
2014-06-25T04:19:06-0400
Let radius is r cm and height is h cm. Then we have the system of equations
r/h=5/7
pi *r2h=550
from the first equation we get h=7r/5. Substitute this into the second equation and we get
pi*7r3/5=550
r=cubic root(2750/4pi)=5 cubic root(22/7pi) approx equals 5.
h=7r/5=7 cubic root(22/7pi) approx equals 7.
Answer: r=5 cubic root(22/7pi) approx equals 5.
h=7 cubic root(22/7pi) approx equals 7.